# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v3

@@ 5.2 Concordant Example
Figure 4: EI operation in a concordant case (patient $7$).
@@
Figure 5: DI operation in a concordant case (patient $7$).
In Fig. 5, we represent the operation of the DI algorithm in the same scenario. Hence, Fig. 5(a) reports the $M=10$ channels showing the most significant connectivity changes, while Fig. 5(b)) reports the desynchronization level for each of the same channels. Comparing Fig. 5(b) with Fig. 4(b), we can observe that the energy ratio represents a less noisy metric than the desynchronization level. During the baseline, the energy ratio of each channel tends to zero, while, for $t_{\text{start}}>0$, most channels present abrupt energy changes, with a peak $E_{x}$ for channels Y4 and Y5. The desynchronization level is characterized by higher variability, and many channels present bursts of reduced connectivity for $t_{\text{start}}<0$. In particular, channels Y4 and Y5 present abnormal connectivity patterns even during the baseline.
@@
Figure 6: EI and DI outcomes in a concordant case (patient $7$).
Hence, in Fig. 6, we report the epileptogenicity levels assigned to each channel by the two algorithms, as well as the brain structure with which each channel is associated. In both cases, the output of the algorithms is normalized within the $[0,1]$ range so that values close to $1$ are associated with a greater influence on the diffusion of epileptic discharge. We can appreciate that the outputs are mostly concordant as both techniques identify the most epileptogenic channels within the posterior collateral sulcus and the fusiform gyrus, which were considered triggers of the patient’s seizures.
5.3 Discordant Example
@@
Figure 7: EI operation in a discordant case (patient $8$).
@@
Figure 8: DI operation in a discordant case (patient $8$).
@@
Figure 9: EI and DI outcomes in a discordant case (patient $8$).
In the following, we consider patient $8$, whose SEEG implant (represented in Fig. 3(b)) explored the left fronto-temporo-parietal, area, and includes $116$ recording sites ($13$ electrodes). We report the $M=10$ channels considered the most epileptogenic according to the two approaches, focusing first on the EI algorithm (Fig. 7) and then on the DI algorithm (Fig. 8). As before, the energy ratio associated with each channel presents low values during the baseline and then increases abruptly during the ictal phase. In particular, the channels associated with electrode I, which crosses the temporal pole, maximize the energy ratio a few seconds after $t_{\text{start}}$.
In Fig. 8, we perform the same analysis for the desynchronization level. In this case, the network connectivity is slightly less noisy than the seizure analyzed in Fig. 5 since the channels do not show prominent variations during the baseline period. This difference may be due to the different network sizes: the SEEG implant of patient $8$ includes $36$ more recording sites than that of patient $7$. In other words, the SEEG implant explores a much more focused epileptogenic network and, thus, leads to a more stable connectivity dynamics.
Focusing on the ictal phase ($t_{\text{start}}>0$), we can appreciate that DI identifies desynchronization phenomena in several structures that, instead, are not marked by epileptogenic in terms of energy ratio. As shown in Fig. 8, channels associated with electrodes B, D, and E (which explore the cortical area adjacent the hyppocampus and collateral sulcus) are characterized by a significant decrease in connectivity a few seconds after the start of the ictal phase. None of these channels (but B1) was included in those marked by EI, which can lead to the definition of a more restricted resection area during epilepsy surgery, with a lower chance of totally suppressing the mechanisms underlying seizure activation.
Fig. 9 compares the epileptogenicity levels assigned to each channel by the algorithms, also reporting the cortical locations of the channels within the SEEG implants. In this case, the output of the algorithm is only partially concordant: EI marked the channels associated with the temporal pole as the most epileptogenic, while DI seems to identify the origin of epileptic discharges in the limbic system. Using EI, the most epileptogenic channel is I1, while the algorithm DI places more emphasis on channel B1, B2, E4 and E5. These heterogeneous results denote the importance of integrating multiple techniques, considering both energy-based and connectivity-based metrics, to improve accuracy in the EZ localization.
5.4 Clinical Discussion
@@
When combined with EI, the DI algorithm allows us to greatly improve the accuracy in EZ localization. From a practical perspective, DI marks as epileptogenic the cortical sites that assume a static phase distribution over time and, thus, are not influenced by any other sites explored through the SEEG implant. The key idea behind our methodology is that epileptogenic signals present an independent behavior that cannot be inhibited by the other cortical structures. This is in line with the most recent findings in the field of EZ detection and confirms that the generation of epileptic seizures is allowed by a sudden reduction in network connectivity [47, 33, 48]. In this vision, the regions surrounding the EZ assume an inhibitory function that prevent epileptic discharges during the interictal phase.
@@
6 Conclusion
In this work, we developed a new computational framework for analyzing SEEG signals, with the final goal of identifying EZ in patients with drug-resistant epilepsy. Our method, named DI, uses PTE to estimate the effective connectivity between the SEEG cortical sites and considers as epileptogenic the SEEG channels that present an abrupt desynchronization in the period immediately preceding the propagation of seizures. To evaluate its clinical utility, we implemented DI on a dataset of $11$ patients and compared its performance in identifying epileptogenic sites against the EI algorithm. The results showed that DI leads to a higher AUC than EI in the EZ identification ($AUC=0.80$ vs $AUC=0.74$), and the best performance ($AUC=0.87$) is obtained when combining the two algorithms as a unique detection tool.

### v4

@@ 5.2. Concordant Example
Figure 4. EI operation in a concordant case (patient $7$).
[…]
[EI operation in a concordant case (patient $7$).]EI operation in a concordant case (patient $7$).
@@
Figure 5. DI operation in a concordant case (patient $7$).
[…]
[DI operation in a concordant case (patient $7$).]DI operation in a concordant case (patient $7$).
@@
Figure 6. EI and DI outcomes in a concordant case (patient $7$).
[…]
[EI and DI outcomes in a concordant case (patient $7$).]EI and DI outcomes in a concordant case (patient $7$).
In the following, we better evaluate the operations of the two algorithms, analyzing two specific seizure events. In the first case, the two algorithms return concordant outcomes and, thus, identify the same channels as the most epileptogenic. In the second case, the algorithms’ results are different, which denotes the importance of integrating different methods for the EZ localization. We first consider patient $7$, whose SEEG implant (represented in Fig. 3(a)) explored the right temporo-parieto-occipital area, and includes $152$ different recording sites ($18$ electrodes). In Fig. 4(a), we report the bipolar representation of the $M=10$ channels associated with the strongest variation in the energy ratio during the 50 seconds preceding the propagation of epileptic discharges. Instead, in Fig. 4(b), we report the energy ratio values associated with each channel.
Looking at Fig. 5(b), we can analyze the operation of the DI algorithm in the same scenario. Hence, Fig. 5(a) reports the $M=10$ channels showing the most significant connectivity changes, while Fig. 5(b)) reports the desynchronization level of each channel. Comparing Fig. 5(b) with Fig. 4(b), we can observe that the energy ratio represents a less noisy metric than the desynchronization level. During the baseline, the energy ratio of each channel tends to zero, while, for $t_{\text{start}}>0$, most channels present abrupt energy changes, with a peak $E_{x}$ for channels Y4 and Y5. The desynchronization level is characterized by higher variability, and many channels present bursts of reduced connectivity for $t_{\text{start}}<0$. In particular, channels Y4 and Y5 present abnormal connectivity patterns even during the baseline.
Hence, in Fig. 6(b), we report the epileptogenicity levels assigned to each channel by the two algorithms, as well as the brain structure with which each channel is associated. In both cases, the output of the algorithms is normalized within the $[0,1]$ range so that values close to $1$ are associated with a greater influence on the diffusion of epileptic discharge. We observe that the outputs are mostly concordant as both techniques identify the most epileptogenic channels within the posterior collateral sulcus and the fusiform gyrus, which were considered triggers of the patient’s seizures.
5.3. Discordant Example
@@
Figure 7. EI operation in a discordant case (patient $8$).
[…]
[EI operation in a discordant case (patient $8$).]EI operation in a discordant case (patient $8$).
@@
Figure 8. DI operation in a discordant case (patient $8$).
[…]
[DI operation in a discordant case (patient $8$).]DI operation in a discordant case (patient $8$).
@@
Figure 9. EI and DI outcomes in a discordant case (patient $8$).
[…]
[EI and DI outcomes in a discordant case (patient $8$).]EI and DI outcomes in a discordant case (patient $8$).
In the following, we consider patient $8$, whose SEEG implant (represented in Fig. 3(b)) explored the left fronto-temporo-parietal, area, and includes $116$ recording sites ($13$ electrodes). We report the $M=10$ channels considered the most epileptogenic according to the two approaches, focusing first on the EI algorithm (Fig. 7(b)) and then on the DI algorithm (Fig. 8(b)). As before, the energy ratio associated with each channel presents low values during the baseline and then increases abruptly during the ictal phase. In particular, the channels associated with electrode I, which crosses the temporal pole, maximize the energy ratio a few seconds after $t_{\text{start}}$.
In Fig. 8(b), we show the results obtained using the desynchronization level as the epileptogenic marker. In this case, the network connectivity is slightly less noisy than the seizure analyzed in Fig. 5(b) since the channels do not show prominent variations during the baseline period. This difference may be due to the different network sizes: the SEEG implant of patient $8$ includes $36$ more recording sites than that of patient $7$. In other words, the SEEG implant explores a much more focused epileptogenic network and, thus, leads to a more stable connectivity dynamics.
Focusing on the ictal phase ($t_{\text{start}}>0$), we can appreciate that DI identifies desynchronization phenomena in several structures that, instead, are not marked by epileptogenic in terms of energy ratio. As shown in Fig. 8(b), channels associated with electrodes B, D, and E (which explore the cortical area adjacent the hyppocampus and collateral sulcus) are characterized by a significant decrease in connectivity a few seconds after the start of the ictal phase. None of these channels (but B1) was included in those marked by EI, which can lead to the definition of a more restricted resection area during epilepsy surgery, with a lower chance of totally suppressing the mechanisms underlying seizure activation.
Fig. 9(b) compares the epileptogenicity levels assigned to each channel by the algorithms, also reporting the cortical locations of the channels within the SEEG implant. In this case, the output of the algorithm is only partially concordant: EI marked the channels associated with the temporal pole as the most epileptogenic, while DI seems to identify the origin of epileptic discharges in the limbic system. Using EI, the most epileptogenic channel is I1, while the algorithm DI places more emphasis on channel B1, B2, E4 and E5. These heterogeneous results denote the importance of integrating multiple techniques, considering both energy-based and connectivity-based metrics, to improve accuracy in the EZ localization.
5.4. Clinical Discussion
@@
When combined with EI, the DI algorithm allows us to greatly improve the accuracy in EZ localization. From a practical perspective, DI marks as epileptogenic the cortical sites that assume a static phase distribution over time and, thus, are not influenced by any other sites explored through the SEEG implant. The key idea behind our methodology is that epileptogenic signals present an independent behavior that cannot be inhibited by the other cortical structures. This is in line with the most recent findings in the field of EZ detection and confirms that the generation of epileptic seizures is allowed by a sudden reduction in network connectivity (Coelli et al., 2019; Jiang et al., 2022; Johnson et al., 2023). In this vision, the regions surrounding the EZ assume an inhibitory function that prevent epileptic discharges during the interictal phase.
@@
6. Conclusion
In this work, we developed a new computational framework for analyzing SEEG signals, with the final goal of identifying EZ in patients with drug-resistant epilepsy. Our method, named DI, uses PTE to estimate the effective connectivity between the SEEG cortical sites and considers as epileptogenic the SEEG channels that present an abrupt desynchronization in the period immediately preceding the propagation of seizures. To evaluate its clinical utility, we implemented DI on a dataset of $11$ patients and compared its performance in identifying epileptogenic sites against the EI algorithm. The results showed that DI leads to a higher AUC than EI in the EZ identification ($AUC=0.81$ vs $AUC=0.74$), and the best performance ($AUC=0.87$) is obtained when combining the two algorithms as a unique detection tool.